OpenStudy (dls):

A particle is projected in gravity such that its x and y cordinates are related s: y=ax-bx^2 then speed of projection=?

OpenStudy (dls):

@Yahoo! @ujjwal

OpenStudy (dls):

@ghazi

OpenStudy (dls):

I tried somethings..

OpenStudy (dls):

\[\LARGE y=xtan \theta-\frac{gx^{2}}{2u^{2}\cos^{2} \theta} \]

OpenStudy (ghazi):

\[\frac{ d }{ dx }(ax-bx^2)=(a-2bx) m/s\]

OpenStudy (dls):

comparing.. a=tan theta and b = rest of the stuff

OpenStudy (dls):

what next

OpenStudy (ghazi):

i thought it would be just differentiation i didnt see its a projectile motion

OpenStudy (dls):

its projectile ! :P its "PROJECTED"

OpenStudy (ghazi):

lol yea yea hold on

OpenStudy (dls):

\[\LARGE x=\tan \theta(1-\frac{X}{R}) \] if we compare with this.. \[\LARGE y=x(a-bx) \] divide by a \[\LARGE y=x(1-\frac{bx}{a})\]

OpenStudy (dls):

R=tan theta?perhaps

OpenStudy (ghazi):

when you have made the comparison i guess you can find velocity too , just from the second part , make comparison and your net velocity will be equal to \[V=\sqrt{V _{H}+V _{V}}\]

OpenStudy (dls):

Options: \[\LARGE \sqrt{\frac{g(a-1)}{2b}} \] \[\LARGE \sqrt{\frac{g(a^{2}+1)}{2b}} \] \[\LARGE \sqrt{\frac{g(a+1)}{2b}}\] \[\LARGE \sqrt{\frac{g(a^{2}-1)}{2b}} \]

OpenStudy (ghazi):

in your comparison you have forgotten g

OpenStudy (dls):

1st one?

OpenStudy (ghazi):

2nd one

OpenStudy (dls):

where is G :S

OpenStudy (dls):

\[\LARGE Y= x \tan \theta ({1-\frac{x}{R})}\]

OpenStudy (dls):

\[\LARGE y=x(a-bx)\]

OpenStudy (ghazi):

the thing that i am saying is \[\tan \theta=a\] and \[b=\frac{ g }{ 2u^2\cos^2 \theta }\] apply the two components of velocity and take out the resultant and i guess you will get option B

OpenStudy (dls):

"apply the two components of velocity and take out the resultant " can u be a bit precise?

OpenStudy (ghazi):

hmm if you are concerned about precision , i am trying me best to explain though i know i am unable to explain but i am tryna hard actually what i am saying is to fine Vh and Vv and take resultant

OpenStudy (dls):

HOW to get Vh and Vv :/ we dont hv anythng

OpenStudy (dls):

@experimentX

OpenStudy (experimentx):

find the vertex of that parabola. using that, find the horizontal component of velocity. and also find the time of flight. then find the range of the parabola. using the time of flight, find the horizontal component of velocity. using these two, find the speed.

OpenStudy (dls):

complex

OpenStudy (dls):

i dont even know how to find the vertex

OpenStudy (experimentx):

there is nothing complicated about it. put x=0, you get y=0, at vertex the slope is zero.

OpenStudy (experimentx):

so dy/dx = 0 .. get the value of x, put that x into above and get the value of y.

OpenStudy (dls):

i cant follow up..im not that advanced :/ not your fault..isn't there an easier method? thanks for helping though!

OpenStudy (dls):

why cant we work out with compring

OpenStudy (experimentx):

there are couple of ways of finding the vertex of parabola.